OpenStudy (anonymous):

The sum of the three consecutive integers is 96. what is the largest integer?

OpenStudy (anonymous):

Ok Let the three consecutive integers be 'x', 'x + 1' and 'x+2' x + x+1 + x+2 = 96 3x + 3 = 96 3x = 96-3 3x = 93 x= 93/3 x = 31 The three consecutive integers are 31, 32 and 33. Check the answer: 31 + 32 + 33 = 96 The answer is correct

OpenStudy (anonymous):

Or let x = be the integer (x-1) + x + (x +1) = 96 3* x = 96 x = 96/3 = 32 so the consecutive numbers are 31, 32, 33 Thus the integer is 32

OpenStudy (anonymous):

let x= smallest integer x+1= second third x+2 all these =96 ^^